OpenStudy (anonymous):

solve for x: (2-e^x)(e^(x))=1

OpenStudy (anonymous):

2- e^x=1/ e^x 1/e^x + e^x-2=0 e^2x - 2e^x+1=0 apply (a-b)^2=a^2+b^2-2ab ( e^x-1)^2=0 e^x-1=0 e^x=1 e^x=e^0 x=0,0

OpenStudy (anonymous):

thankssssss